A tiger leaps with an initial velocity of 55.0 km/h at an angle of 13.0 with respect to the horizontal. What are the components of the tiger’s velocity
1
Expert's answer
2019-09-10T13:26:42-0400

The components of the tiger’s velocity



"v_x=v\cos\theta=55.0\times \cos13.0^{\circ}=53.6\:\rm km/h"

"v_y=v\sin\theta=55.0\times \sin13.0^{\circ}=12.4\:\rm km/h"